### v1

# Retail Central Bank Digital Currency:

Abstract
Nations around the world are conducting research into the design of central bank digital currency (CBDC), a new, digital form of money that would be issued by central banks alongside cash and central bank reserves. Retail CBDC would be used by individuals and businesses as form of money suitable for routine commerce. An important motivating factor in the development of retail CBDC is the decline of the popularity of central bank money for retail purchases and the increasing use of digital money created by the private sector for such purposes. The debate about how retail CBDC would be designed and implemented has led to many proposals, which have sparked considerable debate about business models, regulatory frameworks, and the socio-technical role of money in general. Here, we present a critical analysis of the existing proposals. We examine their motivations and themes, as well as their underlying assumptions. We also offer a reflection of the opportunity that retail CBDC represents and suggest a way forward in furtherance of the public interest.
1 Introduction
The rise of the digital economy means that an ever-increasing share of retail transactions are executed electronically, from e-commerce sales over the Internet to in-person payments done via EMV terminals, contactless cards, and mobile apps. For the past thirty years, we have heard the argument that money is memory, that credit is the consummate form of money, and that cash is only really good for illegal activity and is, at best, an imperfect substitute for credit. And yet, digital currency is gaining salience at the same time that consumers are increasingly relying upon bank money to engage with the economy in general. Cash allows consumers to conduct transactions without being subject to the gaze of state or corporate surveillance, which can be used not only to selectively block transactions but to inform lending, insurance, and employment decisions. Around the world, more businesses than ever are moving their distribution online or refusing cash payments at the point of sale, while at the same time, banks are choosing to close their branches and nudge their depositors to make more of their payments online.
Information can always be copied, so it might be tempting to assume that digital assets cannot be possessed and controlled the way physical assets can. But this view is too simplistic: the way digital assets are used, they can exhibit characteristics of possession and control. Drawing from earlier work by Goodell, Toliver, and Nakib [1], and from published standards, we suggest the following definitions:
  • Possession: property of a relationship between an actor and an asset wherein that actor and no other actor can effect changes to the asset, including its destruction, or transfer the asset to another actor
  • Control: property of a relationship between an actor and an asset wherein that actor and no other actor has the means to specify changes to the asset, including the reassignment of this property itself to another actor, that will be considered legitimate
  • Privacy: “the right of an entity (normally an individual or an organization), acting on its own behalf, to determine the degree to which the confidentiality of their private information is maintained” [2]
We are now at a critical juncture in the history of money: Will people be able to possess and control their own money and decide what to do with it, or will all of their payment choices be subject to the oversight and permission of intermediaries? There is an opportunity to use digital technology to preserve the privacy and property rights of everyday people and to frame the public debate about how this will be done.
A tussle is emerging as a result of the changing nature of how the public uses money and the multiplicity of vested interests and business motivations involved in payments. The design requirements for central bank digital currency (CBDC) are currently under debate. Around the world, various proposals aiming to inform the design of digital currency have emerged. In the United States, the central bank conducted a consultation [3], and Congress commissioned a report [4]. In the United Kingdom, the central bank and treasury conducted a joint consultation [5] following other consultations by the central bank in previous years [6, 7], and Parliament conducted a separate consultation [8]. In the European Union, the central bank conducted a consultation [9], and the European Parliament published a proposed regulation [10]. We shall critically review these proposals in the next section. Other relevant commentary and proposals have been delivered by the central banks of other countries, including but not limited to China [11], Sweden [12], and New Zealand [13], as well as by international organisations such as the G7 [14]. Through this debate, several areas of contention have emerged, including but not limited to the following:
  1. 1.
    Privacy. Consumer privacy is at risk in the digital economy. Many articles have identified the threat posed by a cashless society to the human rights of individual persons, who must engage in the economy by making payments as part of their everyday lives. Indeed, it is trite to say that the purpose of central banks is to ensure safe and equal access of public persons to the economy. Reports acknowledging the threats posed by electronic transactions have been published around the world, by public-sector organisations [15, 16, 17, 18, 19], by private-sector businesses [20, 21], and by civil rights organisations and think tanks across the political spectrum [22, 23]. Central banks have also acknowledged the dangers that cashless payments pose to privacy [24, 25], and the degradation of consumer privacy is also understood to have significant implications for the broader economy [26]. Exceptional access mechanisms, which seek to provide a means by which authorities can gain privileged access to data under certain circumstances, have been consistently dismissed as dangerous and untenable by the security community over the past quarter century [27, 28, 29].
    Data protection by trusted parties is recognised to be an imperfect substitute for not collecting data linking consumers to their transactions in the first instance [30]. Fortunately, alternative approaches are available [31], and the rise of the digital economy does not imply that the relinquishment of privacy by individual consumers is inevitable. In fact, the public utility of privacy in money has been demonstrated [32], and the response to a public consultation on CBDC conducted by the European Central Bank revealed that privacy is the number one concern of respondents [9]. Some designs for CBDC demonstrate that it is possible to offer true and verifiable privacy to consumers [1, 33].
  2. 2.
    Custody. A salient feature of money is the option of individuals to possess and control it directly as a bearer instrument, as they can with cash. However, the custody model described in the relevant proposals in the US [3], UK [5], and Eurozone [10] assume that consumers will not be able to possess or control CBDC. Instead, CBDC would be held in special “wallets”, enabling the use of services offered by providers who would be de facto custodians of the money. The implicit definition of the term “wallet” as a service contrasts with the internationally accepted definition of a wallet as an “application or mechanism to generate, manage, store or use” digital assets [34] (emphasis added).
    Direct custody of money underpins the free exercise of choice in market economies, and if consumers cannot be in direct possession and control of their money, then it is not really theirs. Instead, they would be forced to contend with the possibility that their use of money might be restricted to certain purposes at certain times, as it was for recipients of the controversial cashless welfare card in Australia [35], or curtailed completely, as it was for Canadian lorry drivers in early 2022 [36], who were subsequently exposed to public blacklisting [37].
    Requiring that intermediaries stand between consumers and their own money is tantamount to requiring that consumers have money within custodial accounts. We question whether, from the perspective of a consumer, “digital currency” residing within an account of this type is significantly different from bank deposits residing within ordinary bank accounts. We argue that digital currency must be designed so that, like cash, it can be possessed independently of an accounting relationship.
  3. 3.
    Role of identity. Separately from the question of custody, which is about whether individuals can possess and control their own assets, is the question of the form that the assets take. Digital currency assets are fungible, so in principle they can take the form of either balances (“account-based access”) or tokens (“token-based access”) [38]. Both approaches are technically feasible: Some cryptocurrencies, such as Bitcoin, conduct transactions using tokens, while others, such as Ethereum, conduct transactions using balances.
    However, the choice of balances versus tokens has implications for the role of identity in accessing the assets. In a digital currency system with balances, a number representing the size of some collection assets must be reported somewhere, and successive transactions must result in changes to that number. The state of a balance, therefore, is determined by the net effect of a set of transactions that are associated with that balance and with each other. The persistent linkage among transactions implies the existence of an identifier for the owner of the assets, as well as an implicit relationship between this identifier and all of the transactions made by the owner, as the owner of the assets must provide identification to access the balance to conduct a transaction. Such ownership can be economised but not completely eliminated. Conversely, with tokens, all that is needed to conduct a transaction is knowledge of specific tokens or the cryptographic keys that unlock them. A token is not necessarily connected to other tokens, and successive transactions made by the same person might not be linkable on the ledger. This is what happens with cash, which is, indeed, a kind of token.
    The independence of a user’s transactions is essential for both privacy and control: Whether or not a particular gatekeeper has custody, the requirement to use a balance implies that the balance can be locked as a mechanism for preventing the owner from accessing the assets, or that links among successive transactions can be forcibly discovered as a means of surveillance, as is the case with the development of e-CNY system in China [39].
  4. 4.
    Role of the central bank in processing transactions. Cash payments are naturally decentralised; any two counterparties can consummate a transaction without involving the issuer by exchanging a physical token. Modern electronic payments are also decentralised; they are carried out by networks of private-sector actors overseen by regulators [40, 41].
    Trusting the central bank to operate (or delegate the operation of) a transaction processing infrastructure is fundamentally different from trusting the central bank to oversee a transaction processing infrastructure as a regulator [42]. In particular, if a central actor were to control the system that maintains the records of transactions, then there would be no technical mechanism to stop this central actor from changing the historical record. The best we might expect is for changes to historical data to be observable by periodic reconciliation activities among third parties with read access to the records, followed by a legal challenge wherein the authoritativeness of the records would be determined by indexing the intentions of the maintainer of the records in some juridical context. In the absence of a clear procedure for determining which version of history is true, all parties would be forced to accept the state of transactions decided by an authority, irrespective of what actually transpired. Even if we assume that the authority is not corrupt, all participants in the system must depend upon the integrity of its agents and the effectiveness of its security practices. Recent history teaches us that such dependence is a risky proposition [43].
    Finally, when the party that processes transactions also sets the rules, there is nothing to prevent the rules from changing without warning. Given how many retail consumers would rely upon a future retail CBDC infrastructure, this risk is too much to bear. In distributed models for the operation of best-execution networks, such as the National Market System in the United States [44], the regulator sets the rules but requires participating private-sector actors to operate the system. Changes must be proposed, requested by the regulator, and implemented by all participating actors before they can take effect. This procedural approach protects the system from hazards and is possible only because the system is decentralised in practice. Irrespective of consumer-facing service provision by intermediaries or the potential use of privacy-enhancing technologies, many of the proposed designs for CBDC assume that the central bank or some other centrally operated entity would process individual retail transactions directly.
These areas of contention feature prominently in the various proposals for CBDC that have emerged in recent years, and they demonstrate that the future of money is not simply a technical and economic issue, but fundamentally political in nature. The implementation of retail CBDC in economies with well-established digital payments infrastructure would impact their payment systems at different levels, from the technical underpinning of payments infrastructure to the profitability of consumer banking, from geopolitical aspects of money laundering regulation to the feasibility of delivering economic stimulus packages. Given the scale and scope of changes, it is understandable that experts and non-experts alike might mistake the forest for the trees.
Many of the concerns raised about the design of CBDC reflect an abundance of caution, particularly relating to the use of CBDC to support criminal activity, the potential to undermine the stability of the business models upon which systemically important providers of financial services have come to rely, and threats to the functions of central banks [45]. We shall engage with each of these concerns carefully, with particular attention to the implied trade-offs between different risks. The rapid evolution of technology sometimes means that the choice to do nothing is not be the conservative option that it might otherwise seem to be. Indeed, we argue that CBDC has been billed as important for all the wrong reasons, many of which reflect a fundamental misunderstanding of what CBDC could (or should) potentially achieve, how it would be used, the role of central banking (or what it should be), and the needs of consumers.
The remainder of this article is organised as follows. In Section 2, we offer a review of proposals under consideration in the United States, United Kingdom, and European Union. In Section 3, we identify and contextualise the main themes that feature strongly in the proposals. In Section 4, we explore the motivating factors behind why institutions and governments are spending time and money on CBDC projects. In Section 5, we offer a critical perspective on the state of public discourse on the nature of money, focussing on underlying assumptions about accounts as the linchpin of the relationship between retail payments and institutional power structures. In Section 6, we argue that as a new form of public money, retail CBDC could be a democratising instrument that would promote prosperity, enable competition, bolster innovation, and secure private property rights, but only if it is implemented in a way that ensures that puts owners first. In Section 7, we conclude, summarising the key challenges and offering a path forward.
2 A critical review of current proposals
Next, we consider the design proposals for CBDC that have been developed by various central banks, including the Federal Reserve System, representing the United States; the Bank of England, representing the United Kingdom; and the European Central Bank, representing the Eurozone in particular as well as the European Union more broadly. All of these central banks have championed long-running initiatives to evaluate policy requirements for CBDC and to develop a view on possible approaches to digital currency system design at a technical level. From their published proposals, we shall show that these initiatives have embraced problematic implications for policy and have broadly ignored important technical possibilities. (We note that we have previously reviewed the CBDC proposal of the People’s Bank of China (PBOC), which shares many of the same characteristics [39].) We ask whether these initiatives have served the interests of incumbent actors in preference to the broader public and nascent businesses, and if their constituents, that is, everyday users participating in retail payments, have been abandoned.
2.1 The United States of America: the Federal Reserve
In January 2022, the Federal Reserve System produced a consultation paper, “Money and Payments in the Age of Digital Transformation”, which describes a set of design requirements for central bank digital currency [3]. Technology and system design are not policy-neutral, and although the authors are careful to note that “the paper is not intended to advance a specific policy outcome” [3], the consultation paper nonetheless introduces a set of assumptions, specifications, and technical characteristics that imply a specific set of policy outcomes:
  • Use of accounts. The consultation paper implicitly assumes that individual users of CBDC would not hold it directly. Instead, they would “access CBDC” via accounts with intermediaries. Although never justified, the assumption that CBDC must share its key functional features with bank accounts, rather than cash, is intrinsic to the proposal.
  • Privacy. The consultation paper also suggests that privacy must be limited to accommodate the “transparency necessary to deter criminal activity” [3], leaving open the question of who would decide whether and how an appropriate balance would be achieved, presumably with the assumption that the arbiters would have access to all of the data and decide on the basis of what had been collected. This treatment of privacy is consistent with an article published in 2021 by the G7, which calls for “balance” between “privacy and inclusion” on one hand and “reducing illicit finance” on the other [14].
    Unfortunately, because there is no way for individuals to verify how a set of data is used once it has been collected, privacy ultimately depends upon restricting the set of data that is collected in the first instance. Therefore, depending upon the operational decisions of arbiters is problematic. However, the authors did not justify their assumption that limiting privacy in the interest of “balance” is necessary. They failed to explore the nuances of whether it might be possible to completely protect the privacy of some parties but not others, or of whether it might be possible to collect some data related to all transactions but not all. Privacy and regulatory compliance are and can be complementary, not contradictory, and specifically, it is possible to design an architecture for digital payments that ensures tax and AML compliance for recipients of money without collecting information that can be used to link payers to information about how they spend their money [1]. However, the authors seem to wrongly assume that solutions like this are impossible.
  • Intermediation and custody. The authors state that the initial analysis by the Federal Reserve suggests that a CBDC system should be “intermediated”, with the expectation that the private sector would offer accounts or digital wallet services. Despite the consideration for non-custodial wallets offered by FinCEN [46], the consultation paper assumes that users would have no way to access CBDC except through accounts maintained by intermediaries. Thus, individuals would not be able to possess and control their own money. But there are other models for bearer instruments that are not considered by the authors of the consultation paper, who seem to assume that if a consumer is able to have possession and control of a CBDC asset, then it must be possible:
    1. 1.
      to have peer-to-peer transactions without any involvement of third parties;
    2. 2.
      to have a transaction without reporting the transaction to regulators; and
    3. 3.
      for both counterparties in a transaction to be anonymous.
    In fact, none of these assumptions are true. The consultation paper ignores the possibility that bearer instruments can be combined with transaction infrastructure that directly enforces compliance rules. It has been shown that an architecture with self-validating tokens and oblivious transfers can be built to allow consumers to directly possess the tokens that they use, while also using blind signatures, as proposed by David Chaum [47, 48], to make the payer anonymous. The asset is absolutely a bearer instrument in the sense that the payer holds a fungible asset directly, does not provide identity information during transactions, and is not subject to the rules of a custodian. However, with a system of this type, which is not considered in the consultation paper, it is possible to combine irrevocable anonymity for payers with complete transparency for payees.
  • Role of identity. The authors assume that a requirement for identification implies that transactions would be associated with a single, unique identity. Individuals would have identity-verified wallets, offered subject to on-boarding procedures of the sort that might apply to bank accounts, and that they would only be able to spend CBDC by identifying themselves to the service provider at the time of the transaction. The authors do not consider that it might be possible to identify users at the time that they receive CBDC, as they do when they receive cash from a bank, without requiring them to identify themselves at the point of sale, as is the case with card payments. In effect, the authors assume, without justification, that the transactions of the future would be more similar to card payments, which require consumers to identify themselves, than to cash, which does not have that requirement.
  • Holding limits. The authors consider financial crises in which many users simultaneously seek to switch from bank deposits to CBDC as a way to reduce their exposure to risks associated with private-sector banks, introducing the risk of bank runs and concomitant instability. The authors suggest that a central bank could “address this risk by limiting the total amount of CBDC an end user could hold” [3], although it is not clear that holding limits imposed on wallets would be more effective than limits on CBDC withdrawals in preventing rapid changes in the aggregate volume of CBDC held by individual consumers. Limits on withdrawals could be imposed on individual consumers and businesses, perhaps with different limits for different users, or they could be imposed across the board during crises, perhaps as part of a rapid response mechanism. It should also be noted that limits to the total amount of CBDC in the system could be managed directly by the central bank without monitoring individual transactions or individual holdings at all, as is the case for cash.
We also note that at the time of the report, the Federal Reserve Bank of Boston had already been engaged in Project Hamilton, a collaboration with the MIT Media Lab following a line of technical research and design that had begun in 2016 [49]. Project Hamilton culminated in a software implementation and analysis [50] that was subsequently presented at NSDI 2023 in Boston [51]. The salient features of this design include a UTXO-based token management system wherein tokens are recorded and spent using a distributed database with centralised control. It is not private by design, decision-making and the establishment of truth are entirely centralised, and the database must capture all of the tokens and transactions. However, the authors acknowledge that a token-based architecture is necessary for privacy and that distributed ledger technology could be used to distribute decision-making and control. Although their proposed architecture does not use privacy-enhancing technology to isolate consumer identities from their spending activities, the authors note that such features could be incorporated into the system. They even cite the longstanding work by Chaum, which demonstrates that strong privacy outcomes for payers can be achieved even with centralised architectures [47], as well as his more recent work with the Swiss National Bank on a design for central bank digital currency with strong privacy for payers [48]. However, the Project Hamilton team did not incorporate these features into its design.
In December 2022, a group of nine members of the US House of Representatives led by Representative Emmer of Minnesota penned an open letter to Susan Collins, the President of the Federal Reserve Bank of Boston at the time, questioning the engagement and role of private sector firms in Project Hamilton and the implications of its vision for CBDC on the privacy of individuals [52]. Three weeks later, the Federal Reserve Bank of Boston announced the conclusion of the project [53]. Meanwhile, representatives on the other end of the political spectrum also raised concerns about the potential for poorly-implemented digital currency to have deleterious repercussions for society. Representative Lynch of Massachusetts introduced a bill promoting the development of a privacy-oriented electronic cash system that would explicitly ensure that individuals would be able to spend money anonymously [54]. However, the bill presumed that the security features would depend upon trusted hardware, which is problematic for many reasons [55].
2.2 The United Kingdom: the Bank of England
In March 2023, the Bank of England and HM Treasury published a consultation paper, “The digital pound: a new form of money for households and businesses?” [5], which is the latest document in a series that includes “Central Bank Digital Currency: opportunities, challenges and design” [6] in 2020 and “New forms of digital money” [7] in 2021. The consultation paper concerns the adoption of central bank digital currency (CBDC) for retail use in the United Kingdom. The consultation paper presents a detailed, prescriptive proposal for the design of CBDC and includes strong positions on a variety of contentious points concerning both policy and technology. The positions demonstrate that the authors had made some problematic assumptions, and the resulting system design appears to favour the interests of certain stakeholders in preference to the interests of the general public, particularly retail consumers. We consider several of the most salient issues below:
  1. 1.
    Motivation and process. Overall, although the consultation paper references a variety of specific benefits of CBDC, it is broadly silent about the underlying motivation for considering implementing CBDC in the UK. Many of the arguments presented throughout the consultation paper depend upon accepting certain assumptions, for example that modern electronic payments are well-suited to serve the interests of those who use them, that negate important arguments for researching and developing CBDC, particularly the need to develop an adequate analogue for cash in the digital economy. The lack of clarity about the motivation for developing CBDC undermines the strength of those arguments and raises doubts about their validity.
  2. 2.
    Privacy. The consultation paper states that transactions will not be anonymous because identifying transacting parties is “needed to prevent financial crime” [5]. The privacy model for CBDC reflected in the consultation paper relies upon third parties to collect and safeguard transaction data, and it is assumed that methods would be available by authorities to access such data. Inevitably, although the Bank of England itself might not be able to de-anonymise payers, other parties, including but perhaps not limited to authorities, would be able to do so, and they would be in a position to unilaterally judge whether data protection is appropriate or not.
    The consultation paper repeatedly asserts that a level of privacy similar to that of card payments ought to be sufficient for CBDC, characterising the digital pound as “privacy protected like cards and bank accounts”, “at least as private as current forms of digital money, such as bank accounts”, “on the same basis as currently with other digital payments and bank accounts more generally”, “providing the same privacy as most of the money we use”, and so on [5]. But payments via card networks or bank transfers are not private, as anyone who has seen a bank statement can confirm. Concluding that this level of privacy is acceptable for public infrastructure simply because most transactions have this level of privacy completely ignores the requirements of the transactions that are in the minority, as well as the legitimate concerns of those individuals who accept this level of privacy not as the result of true consent but simply because there is no alternative.
    The consultation paper also suggests business justifications for data collection, arguing that intermediaries can be expected to “use transaction data to improve existing operations or to offer new customer-facing services” and that there is “public appetite for trading personal information for access to products and services” [5]. Whether or not this is true, it is not at all clear that the Bank of England should consider the support and promotion of such business models to fall within its remit.
  3. 3.
    Custody. The consultation paper suggests that payment interface providers (PIPs) would provide wallets, and although PIPs would “never be in possession of end users’ digital pound funds”, all assets would be represented directly on the core ledger, managed and discoverable by core ledger operators or those with the power to compromise or compel access to the information on the ledger. For this reason, end users would not have possession of their own CBDC funds, either.
    Also, in line with the Federal Reserve paper, the Bank of England consultation paper also states that users would face holding limits, citing monetary and financial “stability” and the likelihood of bank runs resulting from a cascade of withdrawals as the justification for imposing limits upon the amount of CBDC that individuals or businesses may possess [5], although their possession of cash is not subject to similar limits. However, protection against bank runs involving CBDC can be realised by imposing restrictions on withdrawals of CBDC from banks, just as it can be realised by imposing restrictions on withdrawals of cash from banks. Applying a limit to withdrawals rather than holdings is more direct and appropriate mechanism to safeguard against bank runs, since it is really the withdrawals by consumers and businesses, not the total holdings by consumers and businesses, that potentially introduce strain on the balance sheets of banks. The choice to impose the restriction on holdings rather than withdrawals is therefore not justified, and an alternative possibility is more apt. We note that there are no limits to how much cash an individual can possess. Were CBDC to become a de facto replacement for cash in the future, would an important right be lost in the process?
    Finally, the consultation paper rejects the idea of bearer instruments, arguing that “a bearer instrument approach […] would lead to completely anonymous payments” [5]. The argument is similar to the argument made in the consultation paper produced by the Federal Reserve, and it is similarly false.
  4. 4.
    Role of identity. The design described in the consultation paper uses balances rather than tokens, and the implications for the rights of asset owners are not explored or justified. It is assumed that wallets would allow users to “[access] digital pounds, make payments, view balances and transaction history” [5]. The functions of wallets as described are not different than the functions of accounts, especially given the fact that “the private sector […] would provide digital pass-through wallets” [5], along with the implicit assumption that owners of digital pounds would be required to identify themselves to PIPs to access their own money.
  5. 5.
    Role of the ledger and scalability. The CBDC design described in the consultation paper makes use of a “core ledger” that records the assets held by individuals. Independently of the privacy considerations and the question of tokens versus balances, the choice to rely upon the core ledger to record assets in this way implies that all transactions must be recorded on the core ledger. This has two problematic consequences for the scalability of the system.
    The first consequence of relying upon the core ledger to store transactions directly is that users must involve the core ledger in every transaction. Because a transaction cannot be consummated until the core ledger is consulted, the operators of the core ledger must be reachable both by the transacting parties (and, if the ledger is distributed, also by each other) during the course of every transaction. Transacting parties must have network connectivity to the core ledger, and they must wait for any system level delays resulting from other transactions taking place at the same time.
    The second consequence of relying upon the core ledger to store transactions directly is that the ledger grows as a function of the number of transactions, which implies that increasing the number of transactions will introduce stress on the mechanism that maintains the core ledger, creating the opportunity for denial of service by adversarial actors. Cryptocurrency systems with this design, such as Ethereum, typically address this vulnerability by introducing a transaction charge (also known as a “gas fee”), which introduces additional economic frictions as well as fairness considerations related to the question of whose transactions have precedence and whether there should be implicit subsidy for users who might otherwise not be able to afford such charges.
    A better design would not rely upon the representation of assets (tokens or balances) directly on the core ledger (distributed or centralised).
  6. 6.
    Role of the issuer. The CBDC design described in the consultation paper specifies an outsized role for the central bank, which not only issues tokens but also processes transactions, sets the rules, and maintains the historical record. It is specified that “the wallet simply passes instructions from the user to the [Bank’s] core ledger”, indicating that all assets are actually managed centrally, and that the “core ledger would be a single piece of infrastructure” [5]. This highly centralised design is unlike any general-purpose payment system that has ever existed before in the United Kingdom. This design implies that the issuer is required to process transactions directly, introducing a tremendous operational, technical, and legal burden upon the central bank.
    Because the system proposed in the consultation paper is highly centralised, its users are exposed to system failure on the part of the central actor, and the central actor is a high-value target for attackers as well.
In March 2022, the Bank of England announced Project Rosalind, a partnership with the Bank for International Settlements (BIS) for designing an API prototype for retail CBDC based upon the architectural insights articulated in its prior consultations [56]. In October 2022, the Bank of England launched a two-week procurement calling for private-sector businesses that had previously registered with its Digital Marketplace to apply to develop a proof-of-concept implementation of CBDC and to undertake research on offline payments [57]. Then, in June 2023, two weeks prior to the extended closing date of the 2023 “digital pound” consultation, the BIS Innovation Hub revealed the names of the Project Rosalind partners, including the API users, which included Mastercard, Amazon, and IDEMIA, as well as the advisers, which included Stripe, Google, and Visa [58], all of whom are active participants in the incumbent identity-linked digital payments system based on custodial accounts.
On 4 July 2023, less than one week after the close of the consultation, the Bank of England updated the Terms of Reference for its CBDC Technology Forum, including a new call for participation and an action plan to establish subgroups to begin design work based upon the specific architecture that it had described in its consultation paper [59]. Six months later, in January 2024, the Bank of England published its response to the consultation, defending the positions that it had articulated in its proposal, once again leaving them broadly intact [60]. The refusal to consider changes to the design was striking, particularly given the manifest concern articulated by the respondents, especially with respect to the rights of users of money in general and privacy in particular.
2.3 The Eurozone: the European Central Bank
In June 2023, the European Commission (EC) published a proposal for a new European regulation on the establishment of a digital euro [10] along with a press release describing a “Single Currency Package […] to support the use of cash and propose a framework for a digital euro” [61].
Similarly to other CBDC proposals, the specification of the rules for the digital euro in proposed regulation implies a specific set of policy outcomes:
  1. 1.
    Privacy. The proposal specifically expresses support for transactions with a similar level of privacy as cash, specifying that this degree of privacy should apply to “offline” payments, whereas “online” payments should have a similar degree of privacy to the electronic payments that are facilitated by banks and card networks today.
    However, given that retail payments in the EU are increasingly online rather than offline, with the value transacted electronically now exceeding €240 trillion per year [64], it is fair to say that the need for privacy is precisely about payments in which at least one party is online, and the proposal by the EC suggests that for transactions with the digital euro that are not “fully-offline”, that is, in which neither party is connected to a network beyond each other, the degree of privacy will match the degree of privacy offered by prevailing electronic transactions facilitated by banks and credit providers, such as card payments. Unfortunately, the privacy offered by such institutions is “privacy by promise”, as those institutions have access to information specifying not only how consumers spend their money, but also metadata, such as location and device information, when they do. It is fair to say that consumers are subject to profiling and surveillance whenever they spend their money in the prevailing electronic transactions referenced by the EC proposal, so to say that consumers would have the same degree of privacy as they would have in such transactions is to say that consumers would not have privacy when they use the digital euro.
  2. 2.
    Motivation. The arguments favouring the protection of cash acceptance and access described in the Single Currency Package are clear: Persons in the Eurozone have a right to use cash, and this right implies both the right to obtain cash (“access” to physical cash, for example from bank branches, ATMs, post offices, and so on) and the right to use cash in their retail payments (“acceptance” of physical cash by retail merchants, in contrast to a more limited concept of legal tender that refers only to the settlement of debts). So, we can conclude that the EC intends to protect cash as a means of payment. However, the attention given to “fully-offline” payments is curious in this context, since the motivation for CBDC is not really about the unsuitability of cash for offline transactions going forward, but instead is precisely about the well-established decline of cash in favour of transactions in which one party is online [62, 63], specifically both e-commerce via the Internet and card and e-money payments at the point of sale. It is not a stretch, therefore, to say that the focus on offline payments is misdirected effort at best, perhaps even disingenuous, leaving the question of the real intention and of which stakeholders are responsible.
  3. 3.
    Certified hardware for “fully-offline” payments. A careful reading shows that the reference to “offline” payments in the proposal for the digital euro really refers only to “fully-offline” transactions wherein both parties are completely disconnected from any third parties. It has been theoretically established that fair exchange between two parties is impossible without the involvement of a trusted third party [65], so in practice, the assumption that “fully-offline” payments are possible would necessitate the involvement of a third-party actor at the point of sale. This is ostensibly achievable if at least one party to the transaction has a device that carries out the will of a third party in preference to the will of either of the transacting parties, although in practice that implies the use of so-called “trusted computing” (also known as a “secure element”) [66, 67] that can attest to a third party that the device had or had not taken a certain action. This kind of remote attestation has been roundly criticised by technical experts and others who advocate for the rights of users exercise control over their own devices [68]. We have previously argued that it would be unreasonable to impose certified hardware requirements on non-custodial wallets [55]:
    1. (a)
      As a basis for the security of a digital money architecture, the security model of trusted computing introduces systemic risk,
    2. (b)
      The inability for a user to analyse the behaviour of a device undermines trust,
    3. (c)
      The inability of experts from the public to produce compatible devices on their own discourages innovation,
    4. (d)
      The special privilege granted to manufacturers of trusted devices is incompatible with market-based competition, and
    5. (e)
      The trusted hardware constitutes a de facto custodian, with a relationship that is likely to carry high switching costs.
    Instead, hardware wallets should focus mainly upon transactions in which a third party is externally reachable, or in which a third party has previously facilitated the transfer of control without possession [1].
  4. 4.
    Use of accounts. Concordantly, the EC proposal makes many references to “digital euro payment accounts” [10], implicitly suggesting that this would be the primary mechanism by which individual persons would access the digital euro. By suggesting that the digital euro is to be “available to natural and legal persons” [10], what the EC proposal seems to suggest is that service providers would offer a new kind of account that would use the digital euro on the back-end to facilitate payments, and that the digital euro would not be held directly by individual owners. And so, the architecture of interbank and card-based payment systems, along with the business model enabled by that architecture, would be preserved in preference to the architecture and business models associated with cash-based payment systems. Providers of “digital euro payment accounts” would have the technical ability to monitor and restrict how a user spends his or her money, just as providers of bank accounts or credit facilities do today.
  5. 5.
    Custody. Even whilst such custodial services might be practical in some cases, self-custody is prerequisite for both privacy and genuine end-user ownership of money. “Wallet” services provided by third-parties are no substitute for non-custodial wallets in which individuals possess and control their own assets. Additionally, the suggested integration with the European Digital Identity Wallet also introduces the possibility that the assets that a natural person owns and how that person spends them would be directly associated with the singular, physical identity of that person. Finally, like the Bank of England and HM Treasury in the UK, the EC proposal also envisions implementing holding limits on the wallets of individuals, rather than considering other ways to prevent individuals from acquiring CBDC too quickly. Considered collectively, the characteristics envisioned by the EC for how digital euro assets are held, if implemented, would have a chilling and manipulative effect on how people spend their money, raising the question of whether their money actually belongs to them or is merely licensed to them on a contingent basis.
The EC proposal for the digital euro does not explain the role that “fully-offline” payments would play in the future economy. If it is assumed that preserving cash is important, then why should the proposal focus so much on “fully-offline” payments, when the existing cash system is both practical and sufficient for this use case? Furthermore, what is the basis for assuming that the privacy of “fully-offline” payments is the only privacy by design worth protecting, when online payments are substituting for cash in practice and are increasingly important to everyday life in Europe? Research suggests that an economy that supports both CBDC and cash might have lower welfare than an economy that supports either form of money alone [69] and that the deployment of a cash-like CBDC that competes with cash might make using cash less efficient [70]. If this turns out to be the case, and “fully-offline” CBDC turns out to be successful, then it is difficult to imagine how the ECB will be able to maintain its headline support of cash for long. More importantly, by focussing the deployment of cash-like CBDC in precisely the environments in which it will compete with cash rather than with custodial payments, the ECB will have missed an opportunity to preserve the most beneficial features of cash, with their concomitant benefits for privacy and self-custody, in the burgeoning digital economy. Instead, the effect would be to bury what remains of those benefits.
3 Themes
As can be seen from the well-publicised digital currency proposals around the world, some common assumptions have begun to materialise. Whether these assumptions result from the influence of incumbent businesses, or the efforts of trans-national regulatory organisations, or the echo chamber resulting from the sharing of ideas by central banks, or other geopolitical factors, their similarities are no coincidence.
3.1 Asset ownership
The report delivered by the Federal Reserve suggests that the litmus test for CBDC should be its added value over existing payment mechanisms, implicitly assuming that existing payment mechanisms will continue to exist and be useful in all relevant contexts. The report articulates a set of assumptions about the design of CBDC and argues that CBDC must offer specific benefits or improvements over existing methods of payment, while offering only a muted acknowledgement of the decline of the use of cash in retail transactions [3]. The report reiterates the US commitment to “ensuring the continued safety and availability of cash” but completely ignores the costs of maintaining cash infrastructure. Such costs have been acknowledged as important by other central banks around the world, including the People’s Bank of China (PBOC), which cited the high fixed costs of cash as a primary motivation for CBDC [11] as well as the central bank of Sweden, which has acknowledged that for cash to survive, it must be explicitly protected [12]. The Reserve Bank of New Zealand echoed this observation, arguing that money requires special “stewardship” that had not been required in the past [13]. If the decline of cash is indeed a driver for the development of retail CBDC, whether for cost or consumer behaviour or other reasons, it is because the properties and value add that cash promotes are worth protecting. Were this not the case, then there would be no value in protecting cash whatsoever.
A report commissioned by the United States Congress and delivered by the Congressional Research Service (CRS) similarly ignores considerations of cash-like properties, which include the ability of a user of cash to directly possess and control it [4]. In its description of the differences between CBDC and the current system, the report seems to suggest that the primary weaknesses of the current system are counterparty risk and costs to users of money, such as personal convenience and system-level efficiency [4], and that those weaknesses constitute the main argument for CBDC. The report also asks where individuals “would be permitted to store and access CBDC” [4], suggesting that the choice would be between financial institutions and the central bank itself, thus setting out a straw man of only two options, as though no other options existed. The report completely ignores the possibility of non-custodial or “unhosted” wallets, which the US Financial Crimes Enforcement Network (FinCEN) had explicitly acknowledged as important in its 2020 consultation. FinCEN even suggested the possibility of establishing specific reporting requirements for transactions involving such wallets [46]. The CRS report does not consider non-custodial wallets, omitting a salient motivation for digital currency in its assessment of all wallets as not fundamentally different from accounts.
These examples highlight, at first instance, the framing of the debate. Instead of taking a holistic approach to the development of money and payments across each of its mediums, it instead looks exclusively at the digital, comparing and contrasting existing electronic retail payment mechanisms that are wholly mediated by custodial accounts with possible CBDC designs, rather than, and more importantly, the much broader landscape of methods and modalities of payment and payment infrastructure, which includes, at its centre, the use of cash and cash-like instruments.
3.2 Privacy and criminal activity
The Federal Reserve report states that privacy and deterrence of criminal activity are fundamentally conflicting values: “Any CBDC would need to strike an appropriate balance, however, between safeguarding the privacy rights of consumers and affording the transparency necessary to deter criminal activity” [3]. The statement that these values are in conflict exposes an assumption that the two desiderata cannot both be satisfied through appropriate design, and it also raises the question of who would adjudicate the “balance”. The report states that CBDC intermediaries must verify the identities of all persons “accessing” CBDC, for the purpose of providing transparency for law enforcement and regulatory compliance [3], similarly to how financial intermediaries maintain records that identify the sender and recipient of transactions between bank accounts. A report published by the UK House of Lords, following a consultation, also assumes that CBDC must track consumer spending behaviour in the same manner as current electronic retail payment mechanisms, which are mediated by custodial accounts, and further suggests that concerns about the lack of consumer privacy might be an inexorable aspect of CBDC in general [8].
However, and contrary to the presumptions of each of these two reports, digital currency differs from bank accounts precisely because it can be stored in unhosted wallets outside custodial relationships, a longstanding feature of cash payments. If the requirement for identity verification is construed to imply that every transaction must be linked to the identities of both of its counterparties, either via the prevailing model for custodial bank accounts or via wallet identification, then consumers would not have privacy when they use CBDC. The assumption that regulatory compliance and crime prevention imply the ability to compromise the privacy of consumers by tracing all payments to their payers is not justified. Although the Federal Reserve report does not state this assumption explicitly, it is implied by the idea that intermediaries could leverage existing tools for managing electronic transfers, wherein all assets are held by custodians [3].
The House of Lords report states that “[w]idespread adoption of any CBDC would depend upon a high level of public trust” concerning privacy [8]. We suggest that the ability for consumers to retain the ability to engage with the digital economy without exposing their transactions to the risk of profiling should be a primary motivation for CBDC. Various architectures that require identification of recipients without linking transactions to payers have been proposed [48, 1], and with non-custodial wallets, it is technically possible to verify the identities of consumers who receive CBDC without allowing them to be linked to their spending habits. Thus, consumer privacy and compliance objectives can be achieved simultaneously, without relying upon “data protection” by a trusted third party and thereby driving a decentralised trust at the centre of digital payments. By technically limiting the amount of data about consumer transactions that are ever received by third parties, privacy by design significantly mitigates the requirement for public trust.
The G7 report from 2021 goes still further than the House of Lords report and states that privacy is in conflict not only with fighting crime but also with the business models upon which CBDC might be designed to depend, stating that privacy by design could “reduce the range of possible business models in a CBDC system” [14]. The interest in protecting “business models” references the possibility that some stakeholders, such as financial intermediaries or platform service providers, might have interests that are precisely at odds with consumer privacy, specifically in harvesting consumer data for business purposes, and that the willingness of such stakeholders to cooperate should be considered a precondition for deploying new infrastructure for retail payments. It seems that maintaining consumer privacy, something that, in the realm of payments, has been enjoyed through the use of cash, is at odds with private sector commercial interests, which wish to track the spending patterns, location, and habits of consumers. The G7 report seems to imply that privacy ought therefore to be curtailed by to protect those private sector commercial interests. This is clearly at odds with the interests of users and consumers.
3.3 Banking versus payment services
Several of the CBDC proposals express concern about disintermediation, the systematic withdrawal of consumer funds from bank accounts into CBDC wallets, as a risk to financial stability. Specifically, they reference the possibility that CBDC would contribute to bank runs in ways that cash would not [8]. The question of whether CBDC is really a substitute for bank deposits arises from the question of why retail consumers use bank accounts in the first instance. CBDC can be designed so that it is useful primarily as a medium of exchange for making retail payments, rather than a long-term store of value [71]. If designed in this manner, it is not really much different from cash and would not necessarily compete with bank deposits as a store of value. Specifically:
  • Not charging interest. It is possible to design CBDC so that, like cash, it does not earn interest. An article published by the IMF argued that the extent to which CBDC competes with bank deposits is largely determined by whether it pays interest [72], a view echoed by the CRS report [4]. Operationally speaking, it would be natural to assume that CBDC assets would not be rehypothecated, as in most cases it is not possible for a central bank to make risky investments [73]. In any event, if CBDC does not earn interest, or even if it earns interest but yields less than commercial bank deposits, justification for the argument that consumers would actively prefer CBDC to bank deposits is unclear.
  • Limiting withdrawals. It is also possible to design CBDC so that withdrawals of CBDC from financial institutions would be operationally restricted in a manner similar to how withdrawals of cash from financial institutions are restricted. Subject to those design characteristics, CBDC would not present a significant risk of large-scale outflow of funds from consumer bank accounts, both because CBDC is an inferior substitute for bank deposits as a store of value over long periods of time and also because withdrawals would be systematically limited. We also note that nation states have imposed limits to bank withdrawals during periods of crisis. For example, in 2015, Greece prevented account holders from withdrawing more than 60 euros per day to protect its banks as its government negotiated the terms of its bailout [74].
With its digital euro proposal, the European Commission explicitly recognised “safeguarding financial stability and financial intermediation” as an objective that should be addressed by limiting the “excessive use of the digital euro as a store of value” [10]. The European Commission also clearly stated that the primary use of the digital euro would be as a means of payment, and that to achieve this objective, it “should not bear interest” [10]. Similarly, the House of Lords report states that the two “main options” for ameliorating the risk of disintermediation is to “disincentivise use” through unattractive interest rates for retail consumers or to “limit the amount of CBDC that can be held or spent” [8]. There is no particular reason to assume that CBDC must pay interest. In fact, if CBDC were designed to be cash-like, then we might assume that it does not pay interest at all, at least in the first instance. Furthermore, notwithstanding the recommendations of some of the CBDC proposals, limits to the amount of CBDC held by consumers can be achieved indirectly, via withdrawal limits and expiration policies instead, and does not require imposing holding limits or other constraints on non-custodial wallets used by consumers, or any other mechanism that would force consumers into unwanted relationships and necessitate revealing or utilising personally identifiable information when they transact.
We conclude that these proposals and reports implicitly assume that consumers value bank accounts primarily as a means for making electronic payments. We question this assumption, although we note that payment services, which generate revenue from data harvesting and transaction fees, constitute a significant source of income for some financial services businesses [75]. If it turns out to be the case that revenue from such services has become essential to the profitability of commercial banks that serve consumers, then we might reasonably conclude that the role of consumer-facing banks has potentially shifted with the decline of cash. In other words, banks have diversified their core business activities away from banking. Such a conclusion would raise the question of whether consumer-facing banks continue to be fit for purpose, if a significant source of their income derives from data harvesting and transaction fees as opposed to interest on lending.
4 Institutional motivations for retail CBDC
Governments and businesses have supported and promoted the development of retail CBDC for a variety of reasons. Chief among them are preservation of monetary and financial sovereignty, protection of existing business models, and institutional security against misbehaviour on the part of users of money. It is fair to say that these motivations are largely reactionary in nature. This section explores these motivations.
4.1 Monetary and financial sovereignty
An important driving force underpinning the pursuit of CBDC involves institutional concerns about emerging risks to monetary and financial sovereignty. Unsurprisingly, such concerns are of particular interest to central banks and governments, who are observing a rapidly changing payments landscape. As retail consumers face a plethora of new ways to make payments, it has become clear that inaction on digital currency is not a particularly safe choice. The set of risks comprises two broad categories: the availability of mechanisms for consumers to make payments with private money, by which we mean cryptocurrencies, privately issued tokens and stablecoins, and e-money in closed-loop payment systems, and the availability of mechanisms for consumers to make payments with foreign money, the currency issued by the central banks of other countries. Both scenarios can undermine the ability of central banks to implement monetary policy and can impact the monetary and financial sovereignty of a nation and its sovereign tender.
4.1.1 Private money
In recent years, several innovations have highlighted the risks of private money creation. Although these innovations are in varying stages of realisation, private-sector actors have potentially tremendous incentives to pursue them:
  • Closed-loop payment systems. According to its 2021 report, the PBOC considers mobile payment systems to be a primary motivation for the development of CBDC [11], and the reason is clear. Platform service providers such as Alibaba and Tencent have developed payment systems that allow users to make payments within a network of platform users, creating a closed loop [76]. Open-loop payment systems, which are popular in most of the world outside China, are effectively fancy wrappers around traditional banking infrastructure, focusing mostly on user experience and convenience factors. Open-loop payments are debited from one bank account and credited to another. In contrast, closed-loop systems allow transactions between platform users without the involvement of regulated financial intermediaries or traditional payment networks [77]. The platform operator generally has a commingled bank account for its users, and the recipient of a payment generally does not receive the funds into a bank account in his or her name. Instead, the platform operator keeps track of the allocation of funds among its clients and users.
    Closed-loop payment systems, such as Alipay and WeChat Pay, have been criticised by the government of China, which adopted “a series of measures to slow the tech companies, enhance the government’s role, and possibly bring payments back into a bank-centric system” [78]. Government regulators might legitimately fear losing the ability to monitor economic transfers that should be eligible for taxation or compliance checks. In the absence of regulation, closed-loop payment platforms have not only the ability to collect data for profit, but also the ability to withhold such data from authorities. It is plain to see why, following the adoption of closed-loop payment platforms by ordinary users as a general-purpose method for remitting payments, the government of China established rules to rein in operators of those platforms [79].
  • Corporate digital currency. Corporations have a variety of reasons for wanting to facilitate payments using money of their own creation, and the mechanisms deployed for such purposes take a variety of forms. Customer loyalty programmes, such as those offered by airlines or coffee shops, allow consumers to accumulate points in exchange for their patronage, and the points can be redeemed for goods or services with partners participating in the loyalty programmes. In some ways, such programmes resemble closed-loop payment systems, although points are generally useable only for consumption and are not exchangeable for fiat currency. Such programmes can be an important source of revenue for their sponsors, particularly if the sponsor sells the right to participate, invests the proceeds, and buys protection against the risk [80].
    In recent years, businesses have explored the possibility of digital currency as a kind of closed-loop payment system on a global scale. In particular, excitement about digital currency projects by Facebook/Meta (Libra, Diem, Zuck Bucks, and so on) and by payment networks such as MasterCard [81, 82] suggest the possibility that consumers and merchants could dispense with fiat currencies entirely, placing their faith in private issuers instead, with, for example, multi-currency stablecoins.
  • Stablecoins and their analogues. We use the term stablecoin to denote a privately-issued token whose value is pegged to the value of another asset, such as a currency issued by a central bank or a commodity asset. The peg is generally maintained by the issuer. Stablecoins can be further divided into two categories: algorithmic stablecoins, whose value is maintained by trading activity (including issuance and redemption) and collateralised stablecoins, which are backed by assets held by the issuer. Stablecoins backed by central-bank reserves have also been proposed [83].
    From the perspective of risk, stablecoins can be compared to assets traded in money markets such as bankers’ acceptances, commercial paper, and so on. However, unlike those assets, stablecoins can be traded via a digital infrastructure outside of regulatory oversight. The ability of ordinary persons in some country to conduct commerce using stablecoins pegged to something other than the currency of that country might pose a threat to monetary sovereignty, particularly for central banks and governments with poor credit or unstable currency [84]. This concern was specifically recognised when Facebook (now Meta) first introduced the now-defunct Libra project [85]. Meanwhile, regulators of countries with strong currencies have expressed concern about the business model underpinning stablecoins, suggesting the possibility of a tussle in the medium term [86].
4.1.2 Foreign money
Some arguments supporting the adoption of digital currency focus upon the use of foreign money in domestic contexts and the risk associated with state actors more generally. These risks are about not only the adoption of central bank digital currency, but the capture of value associated with the digital economy more generally.
  • Currency substitution. Some governments have expressed concern that domestic consumers already prefer foreign currency. Specifically, reduced barriers to international trade concomitant with the rise of the digital economy may favour the proliferation and use of foreign currency in preference to domestic currency in some countries, particularly governments whose national currencies are free-floating but whose condition is such that they cannot avoid secular inflation. For example, consider El Salvador, whose leaders opted to accept Bitcoin as legal tender [87], ostensibly following pressure to take action to avoid perceived “dollarisation”.
  • Regulations to manage the “spillover” risk of foreign CBDC. Access to the CBDC of foreign countries has the potential to “lead to currency substitution and loss of monetary sovereignty in both the issuing and foreign country, which in turn might impede the ability of authorities to achieve their own policy objectives, including monetary and financial stability, and the countering of illicit finance” [14]. An economic model developed by researchers at the European Central Bank suggests that the use of foreign CBDC could amplify the transmission of shocks and increase systemic risk [88]. Arguably, part of the rush by governments to design CBDC systems is related to the impetus to ensure that they will have a voice in the creation of regulatory mechanisms and standards that will underpin the global rules for CBDC, and that they will be ready to protect themselves from the effects of foreign CBDC in the event that it is deployed.
    One can easily imagine an outcome in which tourists visiting a particular destination opt to use the CBDC of their home country, and that, through platforms of payment service providers, it is easily offered and accepted by merchants in the destination country. Were this to become widespread across tourism, trade, and other areas, it could impact a nation’s monetary and financial sovereignty.
  • Opportunity to create a common currency. The flip side of the currency substitution risk is an opportunity for trans-national communities with strong economic ties to create a similarly strong common currency for trade. Whether realistic or not, digital currency could be seen as an inexpensive and politically feasible way to introduce a mechanism for conducting trade without incurring the costs associated with establishing a banking system in the new currency, managing a new physical currency infrastructure, or negotiating new rules with neighbours.
  • The “Sputnik effect”. Some governments perceive a race to unlock economic potential from digital currency, fuelled by the perception of opportunity by private-sector businesses to capitalise on the potential for digital transformation. This race has geopolitical importance, not only because governments can potentially use CBDC to promote innovation by their domestic businesses, but also because their domestic businesses might be able to capture value from new, international value chains associated with technologies and services that will underpin new mechanisms for digital payments.
    Suppose, for example, that a particular nation builds out a suite of digital tools that it and the businesses within its jurisdiction expect to export to businesses and states of other nations. Part of the provision will include the innovations and intellectual property of that nation and its businesses. Possibly, their own CBDC would be linked to it, and possibly not. In either case, however, that nation will be positioned to support and therefore influence the value chain in the digital economies of other nations, towards capturing some of its value.
4.2 Lobbying opportunities for private-sector incumbents
Much of the opportunity for private-sector businesses is related to the potential to extract economic rent, for example, by providing identity services to consumers or by securing the institutional acceptance of digital currency standards that support their business models. These opportunities are shared between incumbent actors in financial services and the technology sector, including both technology platform operators and the innovation ecosystems that they lead.
  • Banks and financial market participants. Over the past several decades, there has been a secular shift toward payments as a significant source of revenue for consumer-facing banks [75]. Possibly, this is related to the downward trajectory of interest rates over the same period [89] and the fact that low interest rates can reduce the margins that consumer-facing banks earn from deposits, thus impacting their profitability [90]. We surmise that commercial banks were incentivised to develop innovative solutions to maintain profitability in their consumer-facing business lines. Although interest rates have risen in response to inflation in recent years, this new development is relatively recent, and the value chains from payments-oriented businesses and infrastructure have remained. Global payments revenues have grown steadily over the longer period, reaching 1.5 trillion dollars per annum, and analysts are projecting that this growth will continue over the next decade [91]. The possibility that CBDC might substitute for bank deposits is seen as a risk factor with implications for financial stability [92]. Although CBDC can be designed to avoid being used as a long-term store of value [71], it is worth considering whether retail consumers establish bank accounts not primarily as a store of value, as banks have traditionally been used by depositors, but instead as a medium of exchange, a relatively new role of banks as payment service providers first and foremost. Concordantly, it might also be worth considering whether the changing business model of consumer-facing banks introduces a new example of systemic risk.
  • Technology service providers. Banks and prevailing platforms for electronic payments also face a reckoning as new technology platforms offer significant competition through better user experience, cheaper fees, ease of access, cross-border interoperability, and better interoperability across the various ecosystems within existing platforms. Consider the rise of Alipay and WeChat Pay in China [93], which provide a wealth of high-assurance personal data that are used to build detailed profiles of their users, which in turn form the basis for credit decisions about those users [94]. For these reasons, social media and e-commerce platform operators are keen to replicate their success in the West [95]. Certainly there is no particular reason that banks must be the ones to provide payments infrastructure; after all, payments infrastructure is more about telecommunications than banking. Specifically, it is more about the underlying technological infrastructure and operating system and associated standards than it is about being able to accept deposits and offer credit.
  • Data harvesters (a.k.a. “Big Tech”). As we have learned from the rise of surveillance capitalism, there is tremendous value in profiling individuals [96]. As it turns out, payments are excellent channels for accumulating data about individuals. Because of AML regulations demanding strong customer authentication as a way to deter criminal activity, they offer links to high-assurance identity information. Because the transactions themselves cost money, mistakes and security breaches are costly, and as a result, the data quality is higher than it otherwise would be. And finally, because nearly everyone must inexorably interact with the economy, consumers cannot avoid making payments. As e-commerce (that is, payments conducted via the Internet) transactions supplant services traditionally supplied by brick-and-mortar shops, and as merchants increasingly refuse cash at the point of sale, electronic payments have captured an increasing share of all retail payments. The data accumulated from consumer payments has been perceived as a way to monitor credit and therefore reduce risks associated with lending and insurance, and is sometimes considered, somewhat dubiously, to be a vector for promoting financial inclusion.
The assignment of current payments industry incumbents to roles in the future digital payments landscape has not yet been resolved. All of the current market participants face uncertainty, and for them, participation in the framing the narrative for the future of money in general, and CBDC in particular, is a way to manage risk. At the same time, it is also worth noting the conspicuous absence of some participants in the contemporary payments industry, most notably those involved in operating the infrastructure that forms the basis of the cash economy, among the global payments conferences and public debates about the future of money.
This absence is a strategic mistake for the providers of cash services as well as a coup for providers of services that support and encourage consumers to use custodial accounts to conduct payments. The cash economy has supply chains that include the businesses that design and manufacture central bank notes and coins, the providers of security technology for the anti-counterfeiting features of physical cash, ATM networks, cash security and transport businesses, manufacturers of cash registers, and merchant-facing banks that take cash deposits, among other businesses.
If CBDC is about providing a public payment option in the form of digital cash, then the development of CBDC could be viewed less as an extension of existing digital payments systems, and more as an adaptation of physical means of payment to the digital economy. In this case, it would stand to reason that businesses involved in these supply chains would be natural partners and instigators of any plan to create digital cash However, it is not clear that those businesses have had much of a voice in public conversations about CBDC to date.
4.3 Governments, institutional incentives, and the Panopticon
Institutions, including governments, have incentives to support the increased collection of data about individuals. Some of these incentives are related to security practices, which can be understood in terms of the classic tussle between security and privacy. Other incentives are political, ranging from interactions with international rule-making organisations to the perennial task of justifying the use of taxpayer money.
  • Crime and tax evasion. Some governments have associated the use of cash with money laundering and corruption [97], arguing that by providing a universally-accepted digital payment option, CBDC is the key to implementing a cashless society. The governor of the central bank of Bangladesh expressed a desire for all retail transactions to be settled by digital or mobile technology by 2027, arguing that eliminating cash would improve efficiency, improve financial inclusion, and reduce crime [98]. In this context, it is easy to understand why government officials in Bangladesh seem to believe that if “a company doesn’t accept digital payments, or if they strongly discourage it, that behaviour itself should be considered suspect” [99]. Similarly, Godwin Emefiele, governor of the Central Bank of Nigeria, argued that a policy of systematically eliminating the use of cash would facilitate the tracking of funds [100]. The e-Naira, a CBDC, was intended to facilitate the transition to a cashless society, although it was ultimately rejected by the public, which responded with protests [101, 102].
  • Financial inclusion. Many CBDC proposals include designs that track the spending habits of ordinary consumers. The International Monetary Fund published an argument citing CBDC as beneficial to financial inclusion because “CBDC generates greater surplus in lending by reducing credit-risk information asymmetry” provided that “is valuable to households as a means of payment or for credit-building” [103]. The mechanism by which information asymmetry is reduced is the ability of lenders to observe patterns in spending activity among consumers. A similar argument was previously made by the Bank for International Settlements, which published an article stating that “the value of data generated by CBDC payments is also clear” as a means of supporting credit and insurance services to “unserved or underserved segments of the population” [104]. Some experts have argued that unlike digital financial innovation offered by private-sector banking services, CBDC can be promoted as a public payment option, and some experts have suggested that CBDC can “only promote financial inclusion and stability when they serve public interest” [105]. Similarly, the Asian Development Bank argued that “financial exclusion may be better addressed by the central banks by using the data on individuals collected through their use of CBDCs” [106].
  • Institutional security. Unsurprisingly, institutions and businesses are motivated to collect intelligence to support their security procedures, including both prevention and response. Identity information provided with payment transactions can be collected as part of a broader effort to institute procedures for anomaly detection that can potentially aid in the prevention of crime and terrorism that could threaten the well-functioning of institutions.
custodialnon-custodialptcard paymentsinterbank transfers ptpossible opportunityfor CBDC ptptcashbearer vouchers digitalphysical
Figure 1: A representation of the universe of payment mechanisms.
5 Accounts all the time
Throughout the debate about the design of CBDC, the question of whether money would be “account-based” or “token-based” has featured prominently. Much of this debate centred around the assumption that token-based money, which can be non-custodial, is either incompatible with a digital payment system or incompatible with the existence or use of bank accounts. This serves as a false dichotomy between digital money and non-custodial money. The use of accounts is not a neutral design choice: Accounts intrinsically introduce at least some degree of accountability of the account-holder to the party maintaining the account. The real issue is about custody: It is possible for money to be both digital and non-custodial (see Figure 1). There are several lines of argumentation that might explain this misunderstanding:
  • Lack of imagination. In the developed world, much of the public is accustomed to an economy intermediated by smartphones, e-commerce platforms, and card payments. The architects of CBDC proposals are certainly familiar with the reality that bank accounts underpin all of their interaction with digital economy today, and it is conceivable that they have not imagined that the future could be any different. Services provided by third parties underpin all interaction with the digital economy, and accounting relationships backed by strong customer authentication are the way that AML/KYC requirements are satisfied. In this context, it might be reasonable to ask: If cash were to be proposed as a means of payment today, then would it be considered too risky to develop and deploy?
  • Data access and control in the name of institutional security. Governments have an operational interest in collecting all possible data about transactions, both ex ante, for the purpose of detecting anomalous behaviour, and ex post, for the purpose of conducting investigations. Data collection by governments is sometimes supported on the grounds that surveillance facilitates the efficiency of investigators and the prevention of crime and terrorism. The justification for such collection can be couched in terms of the paternalistic responsibility that an institution has to people in its milieu. Private-sector organisations and businesses sometimes use a similar argument to support procedures that create profiles of their customers or employees. However, there are significant risks associated with large-scale data collection, including the risk of collateral damage to institutions and individuals alike resulting from data breaches in general, as well as the cybersecurity risks associated with amassing large volumes of marketable personal data in particular.
  • Geopolitical interests. Similarly, governments are keen to pursue data collection in the interest of cooperation with international law enforcement activities, and national governments also have an incentive to use international rule-making organisations to project power beyond their borders. Consider the recommendations of the Financial Action Task Force (FATF), and Recommendation 16 (the “Travel Rule”) in particular, which specifies that both the payer and recipient of electronic transactions must generally be discoverable [107]. A recommendation adopted in 1996 also states: “Countries should further encourage in general the development of modern and secure techniques of money management […] as a means to encourage the replacement of cash transfers” [108]. In 2022, the FATF published a recommendation explicitly stating that the Travel Rule would require financial institutions “to share relevant originator and beneficiary information alongside virtual asset transactions” and characterising “unhosted wallets” as an “emerging risk” [109].
    The FATF recommendations have sparked controversy, particularly in recent years. Evidence suggests that the FATF has been used as a vehicle for projecting geopolitical power, possibly in violation of the United Nations Charter and the Vienna Convention of 1988 [110]. A report by the Royal United Services Institute argued: “Concrete governance does not underpin the FATF’s decisions, even though they may substantially impact countries and their economies. It can therefore be challenging to explain to government ministers and officials how the FATF has the legitimacy necessary to make such decisions” [111].
    In the European Union, there is an additional motivation to pursue CBDC as a means to promote its continued development of its framework for electronic identification, authentication, and trust services, eIDAS. For example, in 2022, a member of the Executive Board of the Deutsche Bundesbank suggested that the government-issued digital identity wallet proposed in the new eIDAS framework (eIDAS 2.0) would be able to hold a digital euro [112]. The eIDAS 2.0 framework has faced public criticism for its implications for personal privacy in the EU [113, 114]. In particular, a group of prominent civil society organizations, academics, and research institutions signed an open letter arguing that the implicit requirement for an individual to use a government-issued wallet that inexorably links together his or her transactions or credentials would “undermine the privacy of EU citizens” and “create an unprecedented risk for every European in their online and offline life” [115].
  • Data subsidy. Another concern is that governments and other providers of infrastructure might prefer to work with businesses that collect and monetise consumer data for profit. In principle, profits from data revenues can be used to offset the costs of building, deploying, and managing the infrastructure. As a result, public services are cheaper if they are run by data harvesters, and infrastructure providers can justify the choice of such vendors by arguing that saving money supports their duty of care to customers or the public. For example, Facebook (now Meta) offers a controversial service called “Free Basics” to provide Internet services to users in developing markets, without requiring them to pay money (specifically, working with local carriers to “zero-rate” the services), but effectively in exchange for their data [116]. Similarly, in November 2023, the data broker Palantir won a bid to operate the data platform for the UK National Health Service [117], affording it an opportunity to extract value from vast volumes of personal information. The implicit subsidy from data revenues can apply at any stage of a project, and is evident from both the eagerness of data analysis firms to engage with infrastructure data sources and the competitiveness of their bids.
  • Cash as an inferior form of money. A line of argumentation has emerged wherein money is cast as a representation of the history of the behaviour of the actors in an economic system; this argument was championed around the end of the last century by Narayana Kocherlakota [118]. In this framing, cash is seen as an imperfect substitute for credit-based approaches that maintain complete records of the histories of individual actors, and the use of credit, in which payers identify themselves, in preference to cash would yield greater economic efficiency. This argument is supported by the perspective that the main benefit of cash is its use in criminal activity [119] presumably because, after all, parties to have nothing to hide would have nothing to fear from using credit. However, the assumption that privacy implies misbehaviour is problematic in general [120], and Charles Kahn and some central bankers used a model to demonstrate that “the value of [cash] money may derive from its supposed imperfection, from the anonymity that it confers” [32] and that removing cash would result in the loss of trade.
  • Surveillance and the Panopticon effect. Surveillance might be seen as a way to control or address abusive behaviour despite the fact that doing so might discourage legitimate but unorthodox patterns of behaviour. Consider the increasing use of technologies by employers to track individuals via their use of personal devices [121], and the increasing use of web security services, such as those provided by Akamai [122], to force customers of businesses that provide online services to reveal their IP addresses, with the stated goal of incentivising good behaviour. Mechanisms to force customers to provide payment card details are no different. Although it is tempting to imagine that people are better-behaved when they are being watched, it is important to consider that the original Panopticon, envisioned as an efficient architecture of a prison with a central tower [123], was as much an instrument of punishment as a tool of compliance [124]. Since there is no way for ordinary individuals to avoid interacting with the economy, do we really mean to create this kind of infrastructure for the general public? And before we assume that the Panopticon would be effective, we must consider that sufficiently wealthy or powerful individuals will find a means to coerce others to conduct transactions at their behest, and potential criminals can expect to find a vast marketplace for stolen identity credentials, including credit cards issued to victims who satisfied AML/KYC criteria, at their disposal [125].
6 Potentially missed opportunities
As nations grapple with the best CBDC design to develop and deploy over the next several years, there is a reckoning concerning which nations will reap the rewards of creating a viable digital payment infrastructure to underpin and support the digital economy. Nations adopting designs that, due to flaws resulting from not prioritising the right properties and underlying principles, may fall behind in the ultimate offering that is provided to users. Ultimately, the categories are simple. There are those that prioritise the protection of business models or sovereignty-based interests in preference to the interests of users, and there are those that treat public access to the digital economy as part of the raison d’etre of a central bank. For central banks in the latter category, those that incorporate properties of cash, such as a token-based form of money that is accessible to all users, privacy-preserving for consumers, non-discriminatory, and compatible with being held and controlled in non-custodial wallets, will be able to reap rewards across a variety of areas.
6.1 Competition and Innovation
One of the central tenets of market efficiency and the avoidance of monopolies, oligopolies, and cartels is the requirement for competition among market participants. Competition would require payment service providers, as well as providers of supporting services, to deliver the best service at the best price for users. If a market is sufficiently competitive, then organisations that fail to be attentive to the interests, demands, and preferences of their users and consumers would not survive for long. Competition also implies the ability for organisations to innovate and provide differentiated value propositions for their customers and suppliers.
One central question is whether a particular direction for the design of digital currency could be expected to nurture greater competition and innovation than any other. Were the models for digital currency based upon custodial accounts to be adopted, then digital currency would be more akin to bank deposits than to cash, and it would be likely that regulated financial institutions, mostly consumer-facing banks, would provide services to users. This approach can be expected to protect the market structure for electronic payments of today, and we do not envisage any major differences. In contrast, were an approach to digital currency based upon tokens and non-custodial wallets to be adopted instead, then digital currency would be more akin to cash. For the purpose of retail payments, not only would banks, incumbent financial institutions, and money services businesses continue to provide essential services, but other kinds of businesses and organisations would have a place in the value chain as well, including those that provide ancillary services for the purposes of facilitating payments, providing hardware, maintaining software, offering better user experiences, and even offering custody for digital assets. The diversity in the range and types of services, as well as the competition and innovation afforded by lower barriers to entry, would enhance the efficiency and value delivered to users and consumers. Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age, are preferable to those that seek to build a new back-end for custodial business models.
6.2 Privacy
Considering the implications of the recent CBDC design proposals for end-user privacy, one might assume that CBDC is inherently bad for privacy and the human rights that depend upon privacy, but an important nuance remains. The digital economy has an increasingly important role in everyday life for an increasing number of people around the world, and in this digital economy, consumers have little or no real privacy. Every time people tap a debit card at an EMV terminal or enter bank account information into a web form to make a payment, their locations, choices, and habits become part of their permanent record. This is happening already, whether central banks deploy CBDC or not.
Institutionally-supported digital currency, including CBDC, offers an opportunity for institutions to protect the privacy of consumers, not with “privacy by promise”, but with real privacy by design using privacy-enhancing technology. Institutions and governments have every reason to do so. After all, accountability of consumers for their transaction history introduces a form of friction that can undermine trade, and market efficiency depends upon allowing consumers to make their own choices, free of the concern that they will be judged for making one choice in preference to another. In addition, supporting privacy in financial transactions can directly reduce some forms of crime, such as domestic abuse [126].
Irrevocable anonymity for payers is not incompatible with regulation, as might be assumed. Removing the ability to determine the payer in a transaction channel does not entail removing the ability to determine the recipient, or allowing transactions to take place outside the view of regulators more generally. In fact, privacy for payers complements regulations ensuring that recipients are accountable for tax obligations, and a system that supports the ability for payers to verify the identity of recipients can reduce fraud without compromising their anonymity. Finally, by requiring recipients of digital currency to satisfy AML/KYC requirements, we can ensure that all transacting parties, even payers, are within the security envelope of the set of identified users, without linking specific users to the transactions in which they served as payer.
6.3 Property
In many cultures, the right to private property is broadly considered to be a foundational human right. Money is no different. English law, for example, includes the idea of a chose in action, which is something whose value in a transaction indexes the legal right of a particular party to transact it, and a chose in possession, which is something whose value in a transaction is intrinsic to holding it.
Cash might be a chose in action in the sense that it represents an obligation of the Bank of England, but it is also a chose in possession of its bearer. Bank deposits are different; from the perspective of the beneficial owner, they are claims on the assets of the bank, a chose in action only. Bank deposits have also replaced cash, in a growing number of countries, as the primary means by which individual persons make retail purchases, raising the question: If users cannot own their own money, then what can they own?
Recently, legal scholars have considered the status of cryptoassets, and there is a growing consensus among scholars of English law that simply treating them as choses in action is insufficient [127, 128]. As it turns out, it is possible to have direct possession and control of digital assets, and there exist technical mechanisms for exchanging digital assets that do not require identity and accounting infrastructure. More saliently, the UK Parliament recently enacted the Electronic Trade Documents Act, which explicitly offers a path to recognising the possession of certain digital trade documents by their bearers, as well as the rivalrous transfer of such documents from one party to another [129]. Such legal reasoning could form the basis for recognising the rights of individual persons to possess and control their own digital assets, including digital cash.
7 The future of money
Governments, central banks, and prominent institutions have taken a variety of approaches to the future of money. These approaches range from enthusiastic to dismissive, from aggressive to passive, and from open to secretive. At the same time, not all stakeholders have a voice in this debate, and not all of the universe of possible CBDC designs are being considered and represented with equanimity and fairness. Amidst the cacophony, some clear themes have emerged. Collectively, these themes form the basis of the approaches taken by many of the nations, their governments, and central banks whose currencies might one day be minted in digital form, many of which reflect an echo chamber of reverberating, unimpressive CBDC design properties that do not place the public and their needs at the centre of the CBDC design debate as the ultimate beneficiaries of this new public digital payment infrastructure.
The debate about the design of digital currency is not only a debate about what form of payments are available to users. Rather, it is broader: The debate about CBDC design is about framing and planning the financial infrastructure that will power the digital economy, first domestically and in a retail context, and then globally. Now is the time for us to decide upon a vision for a new “Internet of value”, powered by payments, akin to the way early forms of cash and banknotes that were created to address, among other things, the need for individuals to buy and sell things without intermediation. We should encourage policymakers to take a step back from the press releases and ask: What are the properties that digital money should have? We suggest the following desiderata for a public digital currency system:
  • Is non-discriminatory by design.
  • Offers privacy by design for payers.
  • Allows consumers to directly possess money and control its ownership.
  • Supports an open architecture without the need for certified hardware.
A similar debate is taking place vis-à-vis the digital economy in general. The debates covered in this article and the motivations of central banks, policy makers and financial services providers are diverse, and yet, they do not centre around the needs and delivering the greatest value to users and consumers. To the contrary, the consequences of the designs proposed by teams in the US, UK, and EU are in fact detrimental to the interests, preferences, protections, and rights of individuals and households that are ultimately the consumers that would be expected to use a CBDC. Specifically, those individuals and households would not possess and control their own money, but instead would merely have a contingent claim on assets possessed and controlled by others. Furthermore, those individuals and households would be accountable to their custodians for how, where, when, and with whom they spend their money. Indeed, a vast chasm continues to grow between the set of institutionally promoted CBDC design proposals and the set of alternative models that must be publicly debated.
On one hand are privacy-thwarting and custodially-mediated designs that force individuals into relationships, wherein every payment transaction that users make is logged and linked to their identities by someone, somewhere, and wherein users are forced to accept the involvement of intermediaries that can control, censure, monitor, and individually restrict their use of money. On the other hand are private, non-custodial designs, for which it is technically impossible to build profiles and restrictions for individual users or to force users to ask permission from an authority when they spend their money. When framed in this way, as an evolutionary and iterative development for money and payments, and as the plumbing for the digital economy, the need for public debate and appreciation becomes apparent.
Part and parcel with this debate is the establishment of groundwork for the desiderata that will govern the local and global retail payment infrastructure in the first instance, and which will extend to other areas of finance and value. The CBDC proposals seem to assume that all digital money must inexorably be held in accounts, or that there is no technically viable alternative to having third parties with institutional capabilities manage the money owned by natural persons. But argument on the basis of this assumption is casuistry, and it would seem to imply that direct ownership of money, and by extension anything of value that can be used as countertrade in an economic transaction, is a thing of the past. We reject such arguments. For us, the best model for digital currency would feature:
  • tokens that can be possessed directly, rather than balances that can only represent money that is possessed and controlled by others;
  • a security model based upon data and compliance restrictions associated with those who receive money in a transaction, not those who originate it, since every originator was, after all, once a receiver;
  • no holding limits, not because individuals should be always be allowed to receive more money, but because they should be able to hold money on their own terms;
  • a scalable operating model that allows transactions to be processed in a distributed manner rather than via a small number of high-value control points; and
  • support for open-source self-custody wallets that offer confidence to individuals and households that their possession and control of money does not depend upon the performance or practices of third-party de facto custodians.
We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements, rather than treating them as easily curtailed pawns that the G7 so casually sets aside in furtherance of a narrow set of commercial interests.
Acknowledgements
The authors acknowledge the UCL Future of Money Initiative and the Systemic Risk Centre of the London School of Economics.

### v2

@@
-Retail Central Bank Digital Currency:
+Retail Central Bank Digital Currency: Motivations, Opportunities, and Mistakes
 Abstract
@@
 If CBDC is about providing a public payment option in the form of digital cash, then the development of CBDC could be viewed less as an extension of existing digital payments systems, and more as an adaptation of physical means of payment to the digital economy.
-In this case, it would stand to reason that businesses involved in these supply chains would be natural partners and instigators of any plan to create digital cash However, it is not clear that those businesses have had much of a voice in public conversations about CBDC to date.
+In this case, it would stand to reason that businesses involved in the supply chains that support cash would be natural partners and instigators of any plan to create digital cash.
+However, it is not clear that those businesses have had much of a voice in public conversations about CBDC to date.
 4.3 Governments, institutional incentives, and the Panopticon
@@
 Identity information provided with payment transactions can be collected as part of a broader effort to institute procedures for anomaly detection that can potentially aid in the prevention of crime and terrorism that could threaten the well-functioning of institutions.
-custodialnon-custodialptcard paymentsinterbank transfers ptpossible opportunityfor CBDC ptptcashbearer vouchers digitalphysical
+custodialnon-custodialcard paymentsinterbank transfers possible opportunityfor CBDC cashbearer vouchers digitalphysical
 Figure 1: A representation of the universe of payment mechanisms.
@@
 Throughout the debate about the design of CBDC, the question of whether money would be “account-based” or “token-based” has featured prominently.
-Much of this debate centred around the assumption that token-based money, which can be non-custodial, is either incompatible with a digital payment system or incompatible with the existence or use of bank accounts.
+Much of this debate has centred around the assumption that token-based money, which can be non-custodial, is either incompatible with a digital payment system or incompatible with the existence or use of bank accounts.
 This serves as a false dichotomy between digital money and non-custodial money.
@@
 And before we assume that the Panopticon would be effective, we must consider that sufficiently wealthy or powerful individuals will find a means to coerce others to conduct transactions at their behest, and potential criminals can expect to find a vast marketplace for stolen identity credentials, including credit cards issued to victims who satisfied AML/KYC criteria, at their disposal [125].
+  • Maintaining the power to exercise unilateral control.
+A final possibility is couched in the twin mechanisms of surveillance and censorship.
+By collecting volumes of personal data and related information about transactions, coupled with building broad profiles of individual persons who transact, service providers can more readily implement censorship of specific individuals and circumstances to meet particular objectives.
+Although the means to determine who and when to censor is not the same as the power to be able to censor a customer, consumer, or employee, the fact that customers are locked into accounts maintained by specific third-party actors means that the ability for customers to access and control their own assets, or to conduct transfers or transactions, can be limited and restricted, at any time and for any reason, by those third-party actors.
+Since custodians (and therefore governments) already enjoy such power over card and interbank payments, relinquishing that power might require an act of great courage.
 6 Potentially missed opportunities
@@
 One central question is whether a particular direction for the design of digital currency could be expected to nurture greater competition and innovation than any other.
-Were the models for digital currency based upon custodial accounts to be adopted, then digital currency would be more akin to bank deposits than to cash, and it would be likely that regulated financial institutions, mostly consumer-facing banks, would provide services to users.
+Were the models for digital currency based upon custodial accounts to be adopted, then digital currency would be more akin to bank deposits than to cash, and it would be likely that regulated financial institutions, mostly consumer-facing banks, would be the only or primary entities to provide services to users, even when there is no technical reason for which this must be the case.
 This approach can be expected to protect the market structure for electronic payments of today, and we do not envisage any major differences.
 In contrast, were an approach to digital currency based upon tokens and non-custodial wallets to be adopted instead, then digital currency would be more akin to cash.
-For the purpose of retail payments, not only would banks, incumbent financial institutions, and money services businesses continue to provide essential services, but other kinds of businesses and organisations would have a place in the value chain as well, including those that provide ancillary services for the purposes of facilitating payments, providing hardware, maintaining software, offering better user experiences, and even offering custody for digital assets.
+For the purpose of retail payments, not only would banks, incumbent financial institutions, and money services businesses continue to provide essential services, but other kinds of businesses and organisations would have a place in the value chain as well, including those that provide ancillary services for the purposes of facilitating payments, providing hardware, maintaining software, offering better user experiences, and even offering custody for digital assets, perhaps coupled with security and management services.
 The diversity in the range and types of services, as well as the competition and innovation afforded by lower barriers to entry, would enhance the efficiency and value delivered to users and consumers.
 Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age, are preferable to those that seek to build a new back-end for custodial business models.
+An approach based on tokens and non-custodial wallets would offer a greater diversity in the kinds and types of organisations that can service the interface between citizens and the digital economy.
+It would also support a way for people to choose between service providers on an as-needed basis, without being locked into account-based relationships that can be used as an instrument of surveillance and control, thus strengthening the potential for competition and thus innovation in servicing retail payments in the economy.
 6.2 Privacy
@@
 Money is no different.
-English law, for example, includes the idea of a chose in action, which is something whose value in a transaction indexes the legal right of a particular party to transact it, and a chose in possession, which is something whose value in a transaction is intrinsic to holding it.
-Cash might be a chose in action in the sense that it represents an obligation of the Bank of England, but it is also a chose in possession of its bearer.
+English law, for example, includes the idea of a chose in action (a debt), which is something whose value in a transaction indexes the legal right of a particular party to transact it, and a chose in possession, which is something whose value in a transaction is intrinsic to holding it.
+Cash might be a chose in action in the sense that it represents an obligation of the Bank of England, but it is also a chose in possession of its bearer as an instrument that can be held, owned, and possessed.
 Bank deposits are different; from the perspective of the beneficial owner, they are claims on the assets of the bank, a chose in action only.
@@
 Such legal reasoning could form the basis for recognising the rights of individual persons to possess and control their own digital assets, including digital cash.
+6.4 Addressing shortcomings of existing retail payments infrastructure
+Card payment platforms dominate the existing digital payments infrastructure for retail payments.
+As public infrastructure, these payment platforms have important shortcomings.
+For example, we know that the existence of merchant fees means that merchants who charge the same price for cash payments as for card payments implicitly allow customers who pay with cash to subsidise those who pay with cards.
+Although the merchant fees might apply only to the card transactions, those fees amount to an economic tax on the transaction and ultimately increase the equilibrium price for consumers.
+Additionally, the benefits from rewards programmes offered by card schemes accrue differently to different classes of consumers.
+As researchers at the Federal Reserve Bank of Boston observed, “[b]ecause credit card spending and rewards are positively correlated with household income, the payment instrument transfer also induces a regressive transfer from low-income to high-income households in general” [130].
+In addition to intrinsic flaws with card payments, there are extrinsic problems as well.
+The largest card platforms, including Visa and Mastercard, are based in the US, much to the chagrin of central banks and payment system regulators in other countries.
+Given the burgeoning systemic indispensability of card payments to many economies around the world, it would be understandable for businesses and governments of countries outside the US to seek strategic autonomy, if for no other reason than to protect themselves from the possibility that the US government might act adversely, for example, to introduce tariffs, or to force card platforms to withdraw from certain markets entirely.
+For all of these reasons, it might be reasonable to ask whether card payments, a twentieth century invention, remain fit for purpose in the twenty-first century and beyond.
+Central bank digital currency might offer a way to rethink the role of card payments in the economy in general, as well as to provide a public payment option for consumers in the digital economy in particular.
 7 The future of money
@@
 When framed in this way, as an evolutionary and iterative development for money and payments, and as the plumbing for the digital economy, the need for public debate and appreciation becomes apparent.
-Part and parcel with this debate is the establishment of groundwork for the desiderata that will govern the local and global retail payment infrastructure in the first instance, and which will extend to other areas of finance and value.
+Part and parcel with this debate is the establishment of groundwork for the desiderata that will govern the local and global retail payment infrastructure in the first instance, and which will extend to other areas of finance and value as they digitise, broaden, and become more integrated.
 The CBDC proposals seem to assume that all digital money must inexorably be held in accounts, or that there is no technically viable alternative to having third parties with institutional capabilities manage the money owned by natural persons.
-But argument on the basis of this assumption is casuistry, and it would seem to imply that direct ownership of money, and by extension anything of value that can be used as countertrade in an economic transaction, is a thing of the past.
+An argument on the basis of this assumption is casuistry, and it would seem to imply that direct ownership of money, and by extension anything of value that can be used as countertrade in an economic transaction, is a thing of the past.
 We reject such arguments.
-For us, the best model for digital currency would feature:
+For us, the best model for digital currency would feature, in addition to our four desiderata, the following fundamental characteristics:
   • tokens that can be possessed directly, rather than balances that can only represent money that is possessed and controlled by others;
@@
   • support for open-source self-custody wallets that offer confidence to individuals and households that their possession and control of money does not depend upon the performance or practices of third-party de facto custodians.
-We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements, rather than treating them as easily curtailed pawns that the G7 so casually sets aside in furtherance of a narrow set of commercial interests.
+We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements and foundational constraints, rather than treating them as easily curtailed pawns that the G7 and others have so casually sets aside in furtherance of a narrow set of interests.
 Acknowledgements

### v3

@@
     However, given that retail payments in the EU are increasingly online rather than offline, with the value transacted electronically now exceeding €240 trillion per year [64], it is fair to say that the need for privacy is precisely about payments in which at least one party is online, and the proposal by the EC suggests that for transactions with the digital euro that are not “fully-offline”, that is, in which neither party is connected to a network beyond each other, the degree of privacy will match the degree of privacy offered by prevailing electronic transactions facilitated by banks and credit providers, such as card payments.
-Unfortunately, the privacy offered by such institutions is “privacy by promise”, as those institutions have access to information specifying not only how consumers spend their money, but also metadata, such as location and device information, when they do.
+Unfortunately, the privacy offered by such institutions is actually a limited form of confidentiality, or privacy by promise, as those institutions have access to information specifying not only how consumers spend their money, but also metadata, such as location and device information, when they do.
 It is fair to say that consumers are subject to profiling and surveillance whenever they spend their money in the prevailing electronic transactions referenced by the EC proposal, so to say that consumers would have the same degree of privacy as they would have in such transactions is to say that consumers would not have privacy when they use the digital euro.
@@
 It has been theoretically established that fair exchange between two parties is impossible without the involvement of a trusted third party [65], so in practice, the assumption that “fully-offline” payments are possible would necessitate the involvement of a third-party actor at the point of sale.
-This is ostensibly achievable if at least one party to the transaction has a device that carries out the will of a third party in preference to the will of either of the transacting parties, although in practice that implies the use of so-called “trusted computing” (also known as a “secure element”) [66, 67] that can attest to a third party that the device had or had not taken a certain action.
-This kind of remote attestation has been roundly criticised by technical experts and others who advocate for the rights of users exercise control over their own devices [68].
+This is ostensibly achievable if at least one party to the transaction has a device that carries out the will of a third party in preference to the will of either of the transacting parties, although in practice that implies the use of so-called “trusted computing”: a “trusted execution environment” or “secure element” that can attest to a third party that the device had or had not taken a certain action [66, 67].
+This kind of remote attestation has been routinely criticised by technical experts and others who advocate for the rights of users to exercise control over their own devices [68].
 We have previously argued that it would be unreasonable to impose certified hardware requirements on non-custodial wallets [55]:
@@
 Instead, the effect would be to bury what remains of those benefits.
+Thematic element Digital dollar Digital pound Digital euro
+Stated motivation improved efficiency and convenience for retail payments economic growth, digital transformation financial stability, monetary sovereignty, strategic autonomy
+Use of accounts users access CBDC via accounts with intermediaries PIPs provide “wallets” that are de facto accounts service providers offer “digital euro payment accounts”
+Privacy privacy is subordinate to a transparency requirement; payers can be de-anonymised payers can be de-anonymised no privacy for “online” payments; limited confidentiality for “offline” payments inside the security envelope of a service provider
+Custody service providers hold funds in accounts wallets operated by PIPs hold funds service providers hold funds in accounts
+Role of identity services used to access accounts would be “identity verified” an “alias service” links pseudonyms to a verified identity “digital euro payment accounts” are subject to identification requirements
+Holding limits holding limits are applied by intermediaries on a per-user basis holding limits are applied to CBDC wallets holding limits are applied to digital euro payment acounts
+Table 1: Common themes observed among the CBDC proposals.
 3 Themes
-As can be seen from the well-publicised digital currency proposals around the world, some common assumptions have begun to materialise.
-Whether these assumptions result from the influence of incumbent businesses, or the efforts of trans-national regulatory organisations, or the echo chamber resulting from the sharing of ideas by central banks, or other geopolitical factors, their similarities are no coincidence.
+As can be seen from the well-publicised digital currency proposals around the world (see Table 1), some common assumptions have begun to materialise, prompting us to seek an explanation.
+We have observed that incumbent businesses have joined CBDC research programmes orchestrated by central banks, thus seizing the opportunity to shape the narrative toward their existing business models.
+In addition, trans-national regulatory organisations, such as the Financial Action Task Force (FATF), have reacted unfavourably to non-custodial and privacy-enhancing electronic payment systems.
+Meanwhile, central banks have used forums, such as the BIS, to share ideas about the design of CBDC systems without much public debate, thus creating an echo chamber.
+Whether the common assumptions about the design of CBDC result from a combination of these or other geopolitical factors, their similarities are no coincidence.
 3.1 Asset ownership
@@
 Similarly, governments are keen to pursue data collection in the interest of cooperation with international law enforcement activities, and national governments also have an incentive to use international rule-making organisations to project power beyond their borders.
-Consider the recommendations of the Financial Action Task Force (FATF), and Recommendation 16 (the “Travel Rule”) in particular, which specifies that both the payer and recipient of electronic transactions must generally be discoverable [107].
+Consider the recommendations of the FATF, and Recommendation 16 (the “Travel Rule”) in particular, which specifies that both the payer and recipient of electronic transactions must generally be discoverable [107].
 A recommendation adopted in 1996 also states: “Countries should further encourage in general the development of modern and secure techniques of money management […] as a means to encourage the replacement of cash transfers” [108].
@@
 The diversity in the range and types of services, as well as the competition and innovation afforded by lower barriers to entry, would enhance the efficiency and value delivered to users and consumers.
-Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age, are preferable to those that seek to build a new back-end for custodial business models.
+Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age are preferable to those that seek to build a new back-end for custodial business models.
 An approach based on tokens and non-custodial wallets would offer a greater diversity in the kinds and types of organisations that can service the interface between citizens and the digital economy.
@@
 Collectively, these themes form the basis of the approaches taken by many of the nations, their governments, and central banks whose currencies might one day be minted in digital form, many of which reflect an echo chamber of reverberating, unimpressive CBDC design properties that do not place the public and their needs at the centre of the CBDC design debate as the ultimate beneficiaries of this new public digital payment infrastructure.
+Breaking through the noise requires building upon the research that exposes the unsoundness in the foundations underpinning the prevailing proposals for the design of CBDC.
+For example, Hyoung-kyu Chey observed the difference between a “state theory of money”, which posits that the foundational role of money is its value as a unit of account, and a “commodity theory of money”, which posits that the foundational role of money is its value as a medium of exchange [131].
+While the prevailing CBDC proposals seem to assume that the primary role of money is to serve as a unit of account, experience with cryptocurrencies suggests that the need for CBDC arises instead from the role of money as a medium of exchange, highlighting a mismatch between the theoretical framework from which account-based CBDC proposals seem to arise and the constraints facing retail consumers as they pay for goods and services in the digital age.
+Similarly, although some of the justifications given for the proposed designs for CBDC reflect fears about the disintermediation of banks, research by Bibi and Canelli suggests that such fears are overwrought [132].
 The debate about the design of digital currency is not only a debate about what form of payments are available to users.
@@
   • support for open-source self-custody wallets that offer confidence to individuals and households that their possession and control of money does not depend upon the performance or practices of third-party de facto custodians.
-We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements and foundational constraints, rather than treating them as easily curtailed pawns that the G7 and others have so casually sets aside in furtherance of a narrow set of interests.
+We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements and foundational constraints, rather than treating them as easily curtailed pawns that the G7 and others have so casually set aside in furtherance of a narrow set of interests.
 Acknowledgements
-The authors acknowledge the UCL Future of Money Initiative and the Systemic Risk Centre of the London School of Economics.
+The authors acknowledge the support and encouragement of anonymous reviewers, anonymous central bankers, the UCL Future of Money Initiative, and the Systemic Risk Centre at the London School of Economics.